if the domain is 1,3,5,7 and the range is 1,2,3,4 what is the rule of the function
Answer by stanbon(75887)

if the domain is 1,3,5,7 and the range is 1,2,3,4 what is the rule of the function
---
Points: (1,1),(3,2),(5,3),(7,4)
---
slope = (2-1)/(3-1) = 1/2
---
y = mx + b
3 = (1/2)5 + b
3 = 2.5 + b
b = 1/2
----
Equation:
y = (1/2)x + (1/2)
=======================
Cheers,
Stan H.
============